# thanks: ORCID: $0$-0003-2020-0803thanks: ORCID: $0$-0003-1233-4174thanks: ORCID: $0$-0002-7208-4464thanks: ORCID: $0$-0003-0665-266Xthanks: ORCID: $0$-0001-7579-8684

### v1

@@ I Introduction
To this end, we have developed public codes (see Mauropieroni/gw_response and peregrine-gw/saqqara) to simulate and analyze LISA data. The simulator features a modular structure for the signal components, the noise components, and the LISA response functions. In particular, compared to earlier work Alvey et al. (2024), we include all three data channels of LISA and account for the full $\sim 3$ years of effective mission duration Colpi et al. (2024); Seoane et al. (2022). In this paper, we are in particular interested in the impact of more realistic noise modeling on the ability to reconstruct SGWB parameters. The base LISA model is a two-parameter noise model parametrizing the amplitude of the test mass (TM) and optical metrology system (OMS) noise assuming a given spectral shape. Being more realistic, however, the noise levels should be expected to vary in time, nominally lying between the current best estimate and the allocated noise budget Colpi et al. (2024). To investigate the impact of non-stationary noise on the ability to search for SGWBs, we cut the data streams into segments of 11.5 days (motivated by the cadence of scheduled satellite operations) where we assume the noise to be well-described by the two-parameter base LISA noise model. We then draw independent fiducial values for the two noise parameters in each segment from Gaussian distributions motivated by the allocated noise budget Colpi et al. (2024).
@@ II.1 Time Delay Interferometry
Notice that the XYZ, and similarly the AET variables, are defined as linear combinations of the single-link variables. As a consequence, the mapping from the single-link variables to any TDI basis can be defined through a 3 by 6 matrix, say $\chi^{I}_{ij}$, where the index $I$ runs over the three variables in the TDI basis and $ij$ labels the 6 single link variables. Analogously, any Power or Cross Spectral Density (PSD/CSD) can be defined by applying a 3 by 3 by 6 by 6 tensor $\Lambda^{I,J}_{ij,lm}\equiv\chi^{I}_{ij}\chi^{J}_{lm}$ on the single-link PSDs and CSDs.
@@ III.1 Case Studies
  • Case 2 (Time-varying Noise): In the second case study, we investigate how time-varying noise across the various data segments impacts the determination of $\alpha$ and $\gamma$. Specifically, the parameter vector $\theta=(\alpha,\gamma,A_{1},P_{1},...)$ that we evaluate the FIM on is now constructed as follows: fix $\alpha=-11$ and $\gamma=0$ as per Case 1, then sample $N_{s}$ realizations of $(A_{\tau}\sim\mathcal{N}(3,0.6),P_{\tau}\sim\mathcal{N}(15,3))$ and populate the parameter vector. For each realization, the corresponding FIM is computed, inverted to obtain the covariance matrix, and the parameter sensitivities and correlations for $\alpha$ and $\gamma$ are read off from the relevant covariance matrix components. This is then repeated for $N_{\mathrm{samples}}=5000$ realizations, generating a distribution over the expected sensitivities.
@@ III.2 Fisher Information Matrix Construction
Figure 1: FIM forecasts for the two-parameter signal model. The pink curves in the corner plot and the inset indicate the result for non-varying noise parameters. The black curves and the black histograms highlight the results for time-dependent noise parameters, sampled according to Sec. II. The fiducial parameters taken here are given by: $\alpha=-11$, $\gamma=0$, $A=3$, $P=15$, with $N_{s}=100$ data segments, $f_{\mathrm{min}}=3\times 10^{-5}\,\mathrm{Hz}$, $f_{\mathrm{max}}=5\times 10^{-1}\,\mathrm{Hz}$, and $T_{s}=11.5\,\mathrm{days}$.
@@ III.3 Fisher results
This leads to a very clear picture regarding the impact of the time varying noise on the sensitivity to SGWB. In particular, by analysing the data on a segment-by-segment basis, the increased sensitivity from the segments where the noise is reduced compared to the average outweighs the reverse impact. As such, we see that the distribution of errors on signal parameters lies below the benchmark value that would arise from either a) noise which maintains a fixed magnitude, or b) analysing the data assuming the noise remains fixed throughout, via some form of averaging process over segments. This motivates developing an analysis strategy that can leverage this improvement in performance induced by the low-noise segments of the data [^5] As an aside, we note that analysing the data allowing for time varying noise when in fact the noise levels are constant does not result in any information loss, i.e., the FIM results for the four parameter model ($\alpha,\gamma,A,P$), averaged over all segments, agree with those found in Case 1.. In the next section, we explore the extent to which SBI is a useful tool for this purpose and demonstrate that we can recover this expected sensitivity distribution in a full analysis of the data.
@@ IV.1 SBI Analysis Strategy
  • Low-dimensional Data Representations: Arguably, one of the key challenges for analysing LISA data in an SBI context is data storage. To train an SBI algorithm requires data $d$ generated according to the implicit data likelihood $p(d|\theta)$ across a range of possible model parameters $\theta$. Breaking down the inference problem as above already helps in this regard because it allows us to store only data for a single segment but use it to train a network capable of analysing all segments. In addition, because SBI is an implicit likelihood technique, we can actually perform any transformation to the data $d\rightarrow f(d)$ and the set of algorithms can (in principle) learn the corrected likelihood for the compressed or transformed data $p(f(d)|\theta)$. Here, we take a similar approach to Flauger et al. (2021); Caprini et al. (2019) and coarse-grain the data across the almost $500,000$ frequency bins. In particular, we first fix the relevant coarse-grained frequencies before directly averaging the data inside the bins defined by this grid. An example of this coarse-graining process is shown in Fig. 2. In the analysis shown below, we coarse-grain the quadratic data down from $\sim 500,000$ frequency bins to 2000similar-toabsent2000$\sim$ 2000∼ 2000777We have also checked explicitly that we can achieve the same posteriors with far fewer bins than this, at least as low as 200. This immediately decreases the training and inference time and reduces data storage requirements.. This has three key benefits as far as SBI is concerned: firstly, we can retain all the necessary information in the coarse-grained representation (as shown by the MCMC comparison); secondly, the data storage requirements are significantly reduced, since we only need to store coarse-grained data for a single segment; and thirdly, the network training is significantly accelerated, since the architecture only has to process data with $\sim 6000$ dimensions, as opposed to the $\sim$1.5 million that one would nominally start with. We describe the network architecture in more detail below, but it is worth noting that this sort of compression on a segment-by-segment basis may not possible in an MCMC context if the exact likelihood for the compressed data $p(f(d)|\theta)$ cannot be written down explicitly [^8] We note, however, that it is possible to write approximate likelihoods for the compressed data. For example, in Flauger et al. (2021) an approximate likelihood is defined for coarse-grained data that are the result of averaging over many time segments. Alternatively, in Baghi et al. (2023), a Wishart likelihood is used to model data averaged across frequency bins..
@@ IV.2 SBI Implementation and Technical Details
SBI Algorithm: In this work, and the corresponding saqqara code, we use the Truncated Marginal Neural Ratio Estimation (TMNRE) algorithm Miller et al. (2021, 2022) [^10] Note this is not the only option, and the core aspects of our proposed framework will remain unchanged regardless of SBI algorithm, up to and including the compression aspect of the network design. Other options include Neural Likelihood Estimation (NLE) Papamakarios and Murray (2016); Alsing et al. (2019) or Neural Posterior Estimation (NPE) Papamakarios and Murray (2016) and their sequential versions.. This is described in detail in Refs. Alvey et al. (2023); Bhardwaj et al. (2023); Miller et al. (2021, 2022), see in particular Ref. Alvey et al. (2024) where the first version of saqqara was released. To recap, however, the key points of the method are:
  • Given simulation data $(x,\theta)$ such that $x\sim p(x|\theta)$, i.e., $x$ is sampled from the data likelihood under $\theta$, one can construct a binary classification problem between joint samples $(x,\theta)\sim p(x,\theta)=p(x|\theta)p(\theta)$ and marginal samples $(x,\theta)\sim p(x)p(\theta)$. Training data for the latter class is easy to generate from simulation data by simply permuting $\theta$ amongst examples or resampling $\theta\sim p(\theta)$. By parametrizing (typically through a neural network) a classifier $f_{\phi}(x,\theta)$ and optimising the binary cross entropy loss,
@@
    with respect to $\phi$, it can be shown that the optimal classifier is given by $f_{\phi}^{\star}(x,\theta)=\log\left[p(x,\theta)/p(x)p(\theta)\right]=\log\left[p(\theta|x)/p(\theta)\right]$. In the expression above, $\sigma(x)=(1+\exp(-x))^{-1}$ is the sigmoid function. In other words, we can directly obtain the posterior-to-prior ratio by optimising this loss. For low-dimensional problems, this can be evaluated on a suitably dense grid of $\theta$, or it can be efficiently sampled, see, e.g., Ref. Anau Montel et al. (2023). If an alternative such as Neural Posterior Estimation Papamakarios and Murray (2016) was used instead, then the density estimator could be directly sampled or evaluated.
@@
With these facts in mind, given coarse-grained quadratic data $D_{\tau}^{I}=f_{\mathrm{CG}}(d_{\tau}^{I\star}d_{\tau}^{I})$ the first step is to normalise across the dataset both $\log D_{\tau}^{I}$ and $D_{\tau}^{I}$ separately, channel-by-channel and frequency-by-frequency, i.e., we essentially whiten the data independently across A, E, T, and frequency. Then, two copies of a standard resnet architecture with $3$ input channels, $64$ hidden features, and $2$ blocks are applied to the normalised versions of both $\log D_{\tau}^{I}$ and $D_{\tau}^{I}$. These networks output a compressed version of the data with shape (3, # of parameters to infer). The second part of the network takes this learned summary statistics and uses the standard ratio estimator implemented in the swyft code Miller et al. (2021, 2022) to estimate the 2-dimensional ratio $p(\alpha,\gamma|s_{\phi}(D_{\tau}^{I}))/p(\alpha,\gamma)$ where $s_{\phi}(D_{\tau}^{I})$ is the learned summary statistic of the input data $D_{\tau}^{I}$. Importantly, both parts of the network are trained simultaneously under the binary cross-entropy loss defined above to obtain a single estimator $f_{\phi}(D_{\tau}^{I},(\alpha,\gamma))=\tilde{f}_{\phi}(s_{\phi}(D_{\tau}^{I}),(\alpha,\gamma))$ which first compresses the data, then computes the marginal posterior. We have tested that moderate changes to this architecture, e.g., increasing the number of residual blocks, or changing the number of hidden features have no impact on the inference quality.
@@ IV.3 Results
Generalisation of strategy: The general strategy employed in this work is to analyze cosmological SGWBs in LISA data segment-by-segment using a pre-trained network to combine individual segments into one informative posterior. Looking forward, we are interested as to how well this strategy can generalise to more complex scenarios. To do so, we identify some challenges and possible solutions. Firstly, the assumption regarding the independence of individual segments could be broken. This would happen for instance if working with data containing long-lived transient signals that exceed segment boundaries, see Sec. II.4 for a more detailed discussion and possible mitigation strategies. Secondly, realistic signal and noise models will likely be more complicated than the simple templates used here. As long as these can be described by parametric models, i.e., they allow for forward modeling, the tools developed here can be straightforwardly generalized. In fact, the modular structure of our code is set up to facilitate this. On the signal side, this may include complicated analytical templates or template data banks. On the noise side, this may involve noise models with many different noise components, whose parameters may vary independently. However, a truly agnostic modeling of either the signal or the noise requires going beyond the analysis strategies presented here, and we plan to return to this in future work. Finally, there is the challenge of data generation and storage. Specifically, to maintain the efficient data generation step described above, we needed to store both the coarse-grained, squared data as well as the cross-terms (totalling 6 different pieces of data per simulation). As the number of modeling components increases, constructing and storing the data this way introduces combinatorially increasing storage requirements. This motivates the further investigation of linear data compression steps, or indeed analysis carried out directly at the linear level, which would then only scale linearly in the number of components. We will explore these avenues in future work and releases of saqqara.
@@
Strengths of SBI Approach: Beyond the applicability of the methods to more complex setups, there are several benefits compared to traditional methods. The first is the inherent ability of SBI to analyze bespoke/flexible data representations. Here, we use that to our advantage by analysing a coarse-grained version of the data. This speeds up inference substantially (by a factor of approximately $10^{3}$ here given by the ratio of the number of initial frequency points to the number of coarse-grained bins), and reduces storage requirements. At the segment-by-segment level, this compression may not be possible with traditional sampling approaches, since it is not possible to explicitly write down the exact likelihood for the coarse-grained data (see footnote 8). The second major benefit is the ability to amortise the analysis over segments. In particular, by training the network to analyze one segment, we then get the posteriors for the other segments “for free”. In a sequential sampling-based approach, however, each segment must be analyzed on the full frequency grid from scratch each time. For $N_{s}=100$ segments, this can lead to a significant inference cost given a single MCMC run on one uncompressed data segment takes $\mathcal{O}(1)$ hour [^11] As mentioned above in footnotes 6 and 9, we are running this on hardware consisting of 18 (shared) Intel(R) Xeon(R) Platinum 8360Y CPUs.. In comparison, for each set of $N_{s}=100$ segments, given a trained network, we can analyze the full mission dataset in $\mathcal{O}(1)$ minute. A final advantage of the SBI approach is that we can directly target the inference results we are most interested in. For example, using TMNRE (although it is possible with other SBI methods) we can directly access the posterior for the signal parameters, marginalised over the noise model. The important point to note here is that the scaling of this method is in principle independent of the parametric complexity of the noise model, whereas classical sampling run-times must scale with the full dimensionality of the problem.
@@
Open Challenges for SBI Approach: Currently, the main challenges facing SBI surround convergence checks and the identification of mismodeling. Specifically, in the SBI context, we are often in the position where inference results are optimal conditional on simulation dataset size, and network architecture. In other words, for a given problem, it is possible that an increase in training dataset size, or a change in network architecture could improve the sensitivity of the method. This is compared to convergence tests such as the Gelman-Rubin criteria for MCMC. In this work here, of course, we made direct comparisons so could benchmark our performance with Fisher analyses and MCMC runs, however, more generally this motivates the further development of these tools for SBI applications. On a related note, some of the same convergence guarantees are not available for SBI techniques - such as the long-term Markov Chain behaviour of MCMC. As such, there is a certain reliance on stable network training and faithful data representations to obtain accurate posteriors. On the other hand, additional cross checks, such as testing coverage, can be a complementary diagnostic for an analysis pipeline. Finally, as the modeling of LISA data becomes more complex and detailed, the identification of systematic effects and any mismodeling, either via the likelihood or the simulator in the SBI context is a crucial step. Again, we would argue that this need motivates the more general development of goodness-of-fit tools and checks as SBI becomes a more widespread analysis technique. In addition, it underlines the need for several different analysis chains, based on different methods and making different modeling choices.

### v2

@@ I Introduction
To this end, we have developed public codes (see Mauropieroni/gw_response and peregrine-gw/saqqara) to simulate and analyze LISA data. The simulator features a modular structure for the signal components, the noise components, and the LISA response functions. In particular, compared to earlier work Alvey et al. (2024), we include all three data channels of LISA and account for the full $\sim 3$ years of effective mission duration Colpi et al. (2024); Seoane et al. (2022). In this paper, we are in particular interested in the impact of more realistic noise modeling on the ability to reconstruct SGWB parameters. The base LISA model is a two-parameter noise model parametrizing the amplitude of the test mass (TM) and optical metrology system (OMS) noise assuming a given spectral shape. Being more realistic, however, the noise levels should be expected to vary in time, nominally lying between the current best estimate and the allocated noise budget Colpi et al. (2024). To investigate the impact of non-stationary noise on the ability to search for SGWBs, we cut the data streams into segments of 11.5 days (motivated by the cadence of scheduled satellite operations, although we note that any segment length could in principle be used here) where we assume the noise to be well-described by the two-parameter base LISA noise model. We then draw independent fiducial values for the two noise parameters in each segment from Gaussian distributions motivated by the allocated noise budget Colpi et al. (2024).
@@ II.1 Time Delay Interferometry
Notice that the XYZ, and similarly the AET variables, are defined as linear combinations of the single-link variables. As a consequence, the mapping from the single-link variables to any TDI basis (including the relevant time delays) can be defined through a 3 by 6 matrix, say $\chi^{I}_{ij}$, where the index $I$ runs over the three variables in the TDI basis and $ij$ labels the 6 single link variables. Analogously, any Power or Cross Spectral Density (PSD/CSD) can be defined by applying a 3 by 3 by 6 by 6 tensor $\Lambda^{I,J}_{ij,lm}\equiv\chi^{I}_{ij}\chi^{J}_{lm}$ on the single-link PSDs and CSDs.
@@ III.1 Case Studies
  • Case 2 (Time-varying Noise): In the second case study, we investigate how time-varying noise across the various data segments impacts the determination of $\alpha$ and $\gamma$. Specifically, the parameter vector $\theta=(\alpha,\gamma,A_{1},P_{1},...)$ that we evaluate the FIM on is now constructed as follows: fix $\alpha=-11$ and $\gamma=0$ as per Case 1, then sample $N_{s}$ realizations of $(A_{\tau}\sim\mathcal{N}(3,0.6),P_{\tau}\sim\mathcal{N}(15,3))$ and populate the parameter vector. For each realization, the corresponding FIM is computed (which involves summing the individual FIMs across segments), inverted to obtain the covariance matrix, and the parameter sensitivities and correlations for $\alpha$ and $\gamma$ are read off from the relevant covariance matrix components. This is then repeated for $N_{\mathrm{samples}}=5000$ realizations, generating a distribution over the expected sensitivities.
@@ III.2 Fisher Information Matrix Construction
Figure 1: Left: FIM forecasts for the two-parameter signal model. The pink curves in the corner plot and the inset indicate the result for non-varying noise parameters. The black curves and the black histograms highlight the results for time-dependent noise parameters, sampled according to Sec. II. The fiducial parameters taken here are given by: $\alpha=-11$, $\gamma=0$, $A=3$, $P=15$, with $N_{s}=100$ data segments, $f_{\mathrm{min}}=3\times 10^{-5}\,\mathrm{Hz}$, $f_{\mathrm{max}}=5\times 10^{-1}\,\mathrm{Hz}$, and $T_{s}=11.5\,\mathrm{days}$. Right: Demonstration that in the case where the noise parameters are genuinely stationary across the dataset, we should expect the same sensitivity on the signal parameters $\alpha$ and $\gamma$ if we analyze the dataset assuming only one pair of noise parameters $(A=3,P=15)$, or allow them to vary across each segment $(A_{\tau}=3,P_{\tau}=15)$ for $\tau=1...N_{s}$. Naturally, the constraints on the parameters covering the full dataset (blue curves/contours) ($A,P$) are tighter than the constraints on the parameters in any one segment $(A_{\tau},P_{\tau})$ (pink curves/contours). In the upper right panel, we show that nonetheless, the constraints and correlations on the signal parameters are identical to the left-hand scenario.
@@ III.3 Fisher results
This leads to a very clear picture regarding the impact of the time varying noise on the sensitivity to SGWB. In particular, by analysing the data on a segment-by-segment basis, the increased sensitivity from the segments where the noise is reduced compared to the average outweighs the reverse impact. As such, we see that the distribution of errors on signal parameters lies below the benchmark value that would arise from either a) noise which maintains a fixed magnitude, or b) analysing the data assuming the noise remains fixed throughout, via some form of averaging process over segments. This motivates developing an analysis strategy that can leverage this improvement in performance induced by the low-noise segments of the data. We note that analysing the data allowing for time varying noise when in fact the noise levels are constant does not result in any information loss, i.e., the FIM results for the four parameter model ($\alpha,\gamma,A,P$), averaged over all segments, agree with those found in Case 1. This is demonstrated in the right-hand panel of Fig. 1. In the next section, we explore the extent to which SBI is a useful tool for this purpose and demonstrate that we can recover this expected sensitivity distribution in a full analysis of the data.
@@ IV.1 SBI Analysis Strategy
  • Low-dimensional Data Representations: Arguably, one of the key challenges for analysing LISA data in an SBI context is data storage. To train an SBI algorithm requires data $d$ generated according to the implicit data likelihood $p(d|\theta)$ across a range of possible model parameters $\theta$. Breaking down the inference problem as above already helps in this regard because it allows us to store only data for a single segment but use it to train a network capable of analysing all segments. In addition, because SBI is an implicit likelihood technique, we can actually perform any transformation to the data $d\rightarrow f(d)$ and the set of algorithms can (in principle) learn the corrected likelihood for the compressed or transformed data $p(f(d)|\theta)$. Here, we take a similar approach to Flauger et al. (2021); Caprini et al. (2019) and coarse-grain the data across the almost $500,000$ frequency bins [^6] The $500,000$ frequency bins comes from the following computation: we have segments of length $T=11.5\,\mathrm{days}$ which means that we have a minimum frequency resolution of $\Delta f=1/T=10^{-6}\,\mathrm{Hz}$. Taking a minimum frequency of $f_{\mathrm{min}}=3\times 10^{-5}\,\mathrm{Hz}$ and $f_{\mathrm{max}}=5\times 10^{-1}\,\mathrm{Hz}$, the total number of frequency bins in the $N=(f_{\mathrm{max}}-f_{\mathrm{min}})/\Delta f=499,970$ i.e. about 500,000.. In particular, we first fix the relevant coarse-grained frequencies before directly averaging the data inside the bins defined by this grid. An example of this coarse-graining process is shown in Fig. 2. In the analysis shown below, we coarse-grain the quadratic data down from $\sim 500,000$ frequency bins to 2000similar-toabsent2000$\sim$ 2000∼ 2000777We have also checked explicitly that we can achieve the same posteriors with far fewer bins than this, at least as low as 200. This immediately decreases the training and inference time and reduces data storage requirements.. This has three key benefits as far as SBI is concerned: firstly, we can retain all the necessary information in the coarse-grained representation (as shown by the MCMC comparison); secondly, the data storage requirements are significantly reduced, since we only need to store coarse-grained data for a single segment; and thirdly, the network training is significantly accelerated, since the architecture only has to process data with $\sim 6000$ dimensions, as opposed to the $\sim$1.5 million that one would nominally start with. We describe the network architecture in more detail below, but it is worth noting that this sort of compression on a segment-by-segment basis may not possible in an MCMC context if the exact likelihood for the compressed data $p(f(d)|\theta)$ cannot be written down explicitly [^8] We note, however, that it is possible to write approximate likelihoods for the compressed data. For example, in Flauger et al. (2021) an approximate likelihood is defined for coarse-grained data that are the result of averaging over many time segments. Alternatively, in Baghi et al. (2023), a Wishart likelihood is used to model data averaged across frequency bins..
@@ IV.2 SBI Implementation and Technical Details
SBI Algorithm: In this work, and the corresponding saqqara code, we use the Truncated Marginal Neural Ratio Estimation (TMNRE) algorithm Miller et al. (2021, 2022) [^10] Note this is not the only option, and the core aspects of our proposed framework will remain unchanged regardless of SBI algorithm, up to and including the compression aspect of the network design. Other options include Neural Likelihood Estimation (NLE) Papamakarios and Murray (2016); Alsing et al. (2019) or Neural Posterior Estimation (NPE) Papamakarios and Murray (2016) and their sequential versions.. This is described in detail in Refs. Alvey et al. (2023); Bhardwaj et al. (2023); Miller et al. (2021, 2022), see in particular Appendix D in Ref. Alvey et al. (2024) where the first version of saqqara was released and a detailed discussion surrounding the algorithm and design choices are presented. To recap, however, the key points of the method are:
  • Given simulation data $(x,\theta)$ such that $x\sim p(x|\theta)$, i.e., $x$ is sampled from the data likelihood under $\theta$, one can construct a binary classification problem between joint samples $(x,\theta)\sim p(x,\theta)=p(x|\theta)p(\theta)$ and marginal samples $(x,\theta)\sim p(x)p(\theta)$. Training data for the latter class is easy to generate from simulation data by simply permuting $\theta$ amongst examples or resampling $\theta\sim p(\theta)$. By parametrizing (typically through a neural network) a classifier $f_{\phi}(x,\theta)$ and optimising the binary cross entropy loss (which is the standard choice for binary classification tasks to assign one of two classes to data examples, see e.g. Miller et al. (2022) for its use in the context of parameter estimation),
@@
    with respect to $\phi$ (the trainable parameters of the neural network $f_{\phi}$), it can be shown that the optimal classifier is given by $f_{\phi}^{\star}(x,\theta)=\log\left[p(x,\theta)/p(x)p(\theta)\right]=\log\left[p(\theta|x)/p(\theta)\right]$. In the expression above, $\sigma(x)=(1+\exp(-x))^{-1}$ is the sigmoid function. In other words, we can directly obtain the posterior-to-prior ratio by optimising this loss. For low-dimensional problems, this can be evaluated on a suitably dense grid of $\theta$, or it can be efficiently sampled, see, e.g., Ref. Anau Montel et al. (2024). If an alternative such as Neural Posterior Estimation Papamakarios and Murray (2016) was used instead, then the density estimator could be directly sampled or evaluated.
@@
With these facts in mind, given coarse-grained quadratic data $D_{\tau}^{I}=f_{\mathrm{CG}}(d_{\tau}^{I\star}d_{\tau}^{I})$ the first step is to normalise across the dataset both $\log D_{\tau}^{I}$ and $D_{\tau}^{I}$ separately, channel-by-channel and frequency-by-frequency, i.e., we essentially whiten the data independently across A, E, T, and frequency. Then, two copies of a standard resnet architecture with $3$ input channels, $64$ hidden features, and $2$ blocks are applied to the normalised versions of both $\log D_{\tau}^{I}$ and $D_{\tau}^{I}$. These networks output a compressed version of the data with shape (3, # of parameters to infer – which here is two, $\alpha$ and $\gamma$). More broadly, the aim of the resnet part of the architecture is to summarise the data into a form (and lower dimensionality) where the binary classification task is simpler to carry out. This is a standard aspect in many SBI algorithms and outputs a “learned summary statistic”, which is in some sense optimised for the classification task. We note that, of course, it is possible to work with some pre-constructed summary, but for complex datasets/simulators, this hand-crafted summary may lose information, so the ability to simultaneously construct a data summary as part of the algorithm is a useful feature. The second part of the network takes this learned summary statistic and uses the standard ratio estimator implemented in the swyft code Miller et al. (2021, 2022) to estimate the 2-dimensional ratio $p(\alpha,\gamma|s_{\phi}(D_{\tau}^{I}))/p(\alpha,\gamma)$ where $s_{\phi}(D_{\tau}^{I})$ is the learned summary statistic of the input data $D_{\tau}^{I}$. Importantly, both parts of the network are trained simultaneously under the binary cross-entropy loss defined above to obtain a single estimator $f_{\phi}(D_{\tau}^{I},(\alpha,\gamma))=\tilde{f}_{\phi}(s_{\phi}(D_{\tau}^{I}),(\alpha,\gamma))$ which first compresses the data, then computes the marginal posterior. We have tested that moderate changes to this architecture, e.g., increasing the number of residual blocks, or changing the number of hidden features have no impact on the inference quality.
@@ IV.3 Results
Generalisation of strategy: The general strategy employed in this work is to analyze cosmological SGWBs in LISA data segment-by-segment using a pre-trained network to combine individual segments into one informative posterior. Looking forward, we are interested as to how well this strategy can generalise to more complex scenarios. To do so, we identify some challenges and possible solutions. Firstly, the assumption regarding the independence of individual segments could be broken. This would happen for instance if working with data containing long-lived transient signals that exceed segment boundaries, see Sec. II.4 for a more detailed discussion and possible mitigation strategies. Secondly, realistic signal and noise models will likely be more complicated than the simple templates used here. As long as these can be described by parametric models, i.e., they allow for forward modeling, the tools developed here can be straightforwardly generalized. In fact, the modular structure of our code is set up to facilitate this. On the signal side, this may include complicated analytical templates or template data banks. On the noise side, this may involve noise models with many different noise components, whose parameters may vary independently. Moving towards more agnostic models implies implementing more flexible templates, typically with a parameter number that varies dynamically while performing the inference. Exploring such scenarios requires some mechanism for model selection (such as reversible jumps in MCMC) to be in-built into the inference pipeline. In principle, SBI already directly allows for model comparison Jeffrey and Wandelt (2024). However, tackling this issue would correspond to simulating data and training a network over several parameter spaces simultaneously. While, conceptually, there are no clear no-goes in performing this procedure, technically, some parts of our training algorithm (mostly related to truncation) would have to be adapted to face this challenge. We plan to return to this in future work. In this work, we have also assumed that the changes in noise parameters occur at discrete times. In reality, it may be the case for some noise components that the change is more adiabatic, occurring slowly over a segment. It is possible to include this directly in the framework just by allowing the noise parameters to vary in time during a segment and then generate the data accordingly. This is because SBI can still learn to marginalize over this non-stationary noise, and the benefits surrounding the fact there is one stationary cosmological component remain. Indeed, one such example of this could be the background of galactic binaries which vary over the year. In this case we could additionally think about conditioning the trained estimator directly on the time at which the segment starts $\tau$, allowing us to leverage our knowledge of the annual modulation, but maintain the benefits of only having to train one network. Finally, there is the challenge of data generation and storage. Specifically, to maintain the efficient data generation step described above, we needed to store both the coarse-grained, squared data as well as the cross-terms (totalling 6 different pieces of data per simulation). As the number of modeling components increases, constructing and storing the data this way introduces combinatorially increasing storage requirements. This motivates the further investigation of linear data compression steps, or indeed analysis carried out directly at the linear level, which would then only scale linearly in the number of components. This would also allow us to trivially move beyond the (implicit) assumption of a diagonal noise basis, which here is the AET basis. We will explore these avenues in future work and releases of saqqara.
@@
Strengths of SBI Approach: Beyond the applicability of the methods to more complex setups, there are several benefits compared to traditional methods. The first is the inherent ability of SBI to analyze bespoke/flexible data representations. Here, we use that to our advantage by analysing a coarse-grained version of the data. This speeds up inference substantially (by a factor of approximately $10^{3}$ here given by the ratio of the number of initial frequency points to the number of coarse-grained bins), and reduces storage requirements. At the segment-by-segment level, this compression may not be possible with traditional sampling approaches, since it is not possible to explicitly write down the exact likelihood for the coarse-grained data (see footnote 8). The second major benefit is the ability to amortise the analysis over segments. In particular, by training the network to analyze one segment, we then get the posteriors for the other segments “for free”. In a sequential sampling-based approach, however, each segment must be analyzed on the full frequency grid from scratch each time. For $N_{s}=100$ segments, this can lead to a significant inference cost given a single MCMC run on one uncompressed data segment takes $\mathcal{O}(1)$ hour [^11] As mentioned above in footnotes 5 and 9, we are running this on hardware consisting of 18 (shared) Intel(R) Xeon(R) Platinum 8360Y CPUs.. In comparison, for each set of $N_{s}=100$ segments, given a trained network, we can analyze the full mission dataset in $\mathcal{O}(1)$ minute. A final advantage of the SBI approach is that we can directly target the inference results we are most interested in. For example, using TMNRE (although it is possible with other SBI methods) we can directly access the posterior for the signal parameters, marginalised over the noise model. The important point to note here is that the scaling of this method is in principle independent of the parametric complexity of the noise model, whereas classical sampling run-times must scale with the full dimensionality of the problem.
@@
Open Challenges for SBI Approach: Currently, the main challenges facing SBI surround convergence checks and the identification of mismodeling Anau Montel et al. (2025). Specifically, in the SBI context, we are often in the position where inference results are optimal conditional on simulation dataset size, and network architecture. In other words, for a given problem, it is possible that an increase in training dataset size, or a change in network architecture could improve the sensitivity of the method. This is compared to convergence tests such as the Gelman-Rubin criteria for MCMC. In this work here, of course, we made direct comparisons so could benchmark our performance with Fisher analyses and MCMC runs, however, more generally this motivates the further development of these tools for SBI applications. On a related note, some of the same convergence guarantees are not available for SBI techniques - such as the long-term Markov Chain behaviour of MCMC. As such, there is a certain reliance on stable network training and faithful data representations to obtain accurate posteriors. On the other hand, additional cross checks, such as testing coverage, can be a complementary diagnostic for an analysis pipeline. Finally, as the modeling of LISA data becomes more complex and detailed, the identification of systematic effects and any mismodeling, either via the likelihood or the simulator in the SBI context is a crucial step. Again, we would argue that this need motivates the more general development of goodness-of-fit tools and checks as SBI becomes a more widespread analysis technique. In addition, it underlines the need for several different analysis chains, based on different methods and making different modeling choices.